A stone has a weight of 5N in air and 4.5 N in water.What is the mass of the stone?


1
Expert's answer
2021-01-26T17:48:17-0500

By the definition of the weight, we have:


"W=mg,""m=\dfrac{W}{g}=\dfrac{5\ N}{9.8\ \dfrac{m}{s}}=0.51\ kg."